OpenStudy (anonymous):

I need help with a distance/rate/time question. Lisa leaves her house for a road trip and drives at a rate of 50 mph. Keisha leaves her house 1 hour later and drives along the same route at a rate of 60 mph. At any given time t, the distance traveled can be calculated using the formula d = rt, where d represents distance and r represents rate. Which system of equations can be used to find how long after Lisa departs her house she is passed by Keisha?

OpenStudy (dtan5457):

do we make the system or is it given as multiple choice?

OpenStudy (anonymous):

d=50t d=60t-1 d=50t d=60(t-1) d=50t d=59t d=50(t-1) d=60t @dtan5457

OpenStudy (dtan5457):

Let t represent time you know that keisha is going 60 mph, but an hour later. which of these represent that?

OpenStudy (dtan5457):

50t does not change.

OpenStudy (dtan5457):

@GodGirl360

OpenStudy (anonymous):

60(t-1)

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

Yep.

OpenStudy (dtan5457):

d=50t d=60(t-1) you can even test this. 50t=60(t-1) 50t=60t-60 t=6 50 times 6=300 60(6-1)=300 there are on the same level at this point, which keisha will pass any second from the 6 hour mark

OpenStudy (anonymous):

Since that equation is only in B, that will be the answer. d=50t - Lisa drove 50mph for t hours. d=60(t-1) - Keisha drove 60mph for t hours an hour after Lisa drove.

OpenStudy (dtan5457):

yes. and 6 will be the numerical answer

OpenStudy (anonymous):

Ok, Thank you.

OpenStudy (dtan5457):

np :)